how to solve factoring method following equation
3x(x-2)=x(x+3)

Answer by checkley77(12844)

3x(x-2)=x(x+3)
3x^2-6x=x^2+3x
3x^2-x^2-6x-3x=0
2x^2-9x=0
x(2x-9)=0
x=0 ans.
2x-9=0
2x=9
x=9/2
x=4.5 ans.
Proof:
3*4.5(4.5-2)=4.5(4.5+3)
13.5*2.5=4.5*7.5
33.75=33.75